OpenStudy (shamil98):

In 1991, the cost of a first-class stamp was 29 cents, and the inflation rate was 4.6%. If the inflation rate stayed constant, the function C(t) = 0.29(1.046)^t would represent the cost of a stamp as a function of years since 1991. According to this function, after how many years would the cost of a stamp reach 60 cents? Can someone assist me with this problem?

OpenStudy (shamil98):

Do i put this as C(t) = 0.60(1.046)^t and solve for t?..

OpenStudy (therealmeeeee):

its C 16 years

OpenStudy (shamil98):

I don't care for some google-researched answer @TheRealMeeeee , that diminishes the true reason for my question.. which is to learn..

OpenStudy (therealmeeeee):

well its right

OpenStudy (agent0smith):

The cost is 60c. C is the cost, so plug in 0.60 for C. Solve for t.

OpenStudy (shamil98):

So. 0.60 = 0.29 (1.046)^t then?

OpenStudy (therealmeeeee):

and i didnt google it FYI

OpenStudy (agent0smith):

Yes. First divide by 0.29 on both sides, then take logs of both sides.

OpenStudy (shamil98):

Yeah, sure, I didn't put any answer choices LOL so where you did get the " its C 16 years " then?

OpenStudy (shamil98):

Alright thank you @agent0smith

OpenStudy (therealmeeeee):

i worked it out so shut your mouth

OpenStudy (shamil98):

0.60 = 0.29 (1.046)^t 2.06896551724 = 1.046^t log 2.06896551724 / 1.046 = t ?

OpenStudy (shamil98):

correction* log 2.06896551724 / log 1.046 = t t = 16.1662

OpenStudy (therealmeeeee):

stop saying my name dude

OpenStudy (agent0smith):

Yep. Looks right.

OpenStudy (shamil98):

Thanks.